# Counting primes with a given primitive root, uniformly

### v1

@@ Abstract.
The celebrated Artin conjecture on primitive roots asserts that given any integer $g$ which is neither $-1$ nor a perfect square, there is an explicit constant $A(g)>0$ such that the number $\Pi(x;g)$ of primes $p\leq x$ for which $g$ is a primitive root is asymptotically $A(g)\pi(x)$ as $x\to\infty$, where $\pi(x)$ counts the number of primes not exceeding $x$. Artin’s conjecture has remained unsolved since its formulation 98 years ago. Nevertheless, Hooley demonstrated in 1967 that Artin’s conjecture is a consequence of the Generalized Riemann Hypothesis (GRH) for Dedekind zeta functions of certain Kummer extensions over $\mathbb{Q}$. In this paper, we establish the Artin–Hooley asymptotic formula, under GRH, whenever $\log{x}/\log\log{2|g|}\to\infty$. Under GRH, we also show that the least prime $p_{g}$ possessing $g$ as a primitive root satisfies the upper bound $p_{g}=O(\log^{19}(2|g|))$ uniformly for all non-square $g\neq-1$. We conclude with an application to the average value of $p_{g}$ as well as discussion of an analogue concerning the least “almost-primitive” root $p_{g}^{\ast}$.
@@ 1. Introduction
$\Pi(x;g)\sim A(g)\pi(x),\quad\text{as}\quad x\to\infty,$ (1)
for an explicitly given $A(g)>0$.
@@
$A_{0}(g)=\prod_{q\mid h}\bigg{(}1-\frac{1}{q-1}\bigg{)}\prod_{q\nmid h}\bigg{(}1-\frac{1}{q(q-1)}\bigg{)}.$ (2)
@@
$A_{1}(g)=1-\mu(|g_{1}|)\prod_{\begin{subarray}{c}q\mid h\\ q\mid g_{1}\end{subarray}}\frac{1}{q-2}\prod_{\begin{subarray}{c}q\nmid h\\ q\mid g_{1}\end{subarray}}\frac{1}{q^{2}-q-1};$ (3)
@@
It is this value of $A(g)$ for which Artin predicts (1). [^1] Artin’s original 1927 formulation was missing the factor of $A_{1}(g)$. Artin realized the need for $A_{1}(g)$ after learning of computations carried out by the Lehmers. See Stevenhagen’s discussion in [16].
Artin’s conjecture remains unresolved. In fact, to this day there is not a single value of $g$ for which we can show even the weaker assertion that $\Pi(x;g)\to\infty$. (However, work of Heath-Brown [7] implies this holds for at least one of $g=2$, $3$, or $5$.) The most important progress in this direction is a 1967 theorem of Hooley [8], asserting that the full asymptotic relation (1) follows from the Generalized Riemann Hypothesis (GRH). [^2] Here and below, GRH means the Riemann Hypothesis for all Dedekind zeta functions.
@@
The asymptotic formula $\pi(x;g)\sim A(g)\Pi(x;g)$ holds whenever $\log{x}/\log\log{2|g|}\to\infty$. More precisely, there is an absolute constant $x_{0}>0$ for which the following holds: If $g\in\mathcal{G}$ and $x\geq\max\{x_{0},\log^{3}(2|g|)\}$, then
$\Pi(x;g)=A(g)\pi(x)\left(1+O\left(\frac{\log\log{x}}{\log{x}}+\frac{\log\log{2|g|}}{\log{x}}\right)\right).$ (4)
@@
$p_{g}\ll\log^{B}({2|g|}),$ (5)
for a certain absolute constant $B$. Indeed, if $K$ is an admissible value of the implied constant in (4), then (5) holds for any $B>K$. In our next theorem, we pinpoint a numerically explicit value of $B$.
@@
The upper bound (5) holds with $B=19$.
Usually $p_{g}$ is quite small. For instance, $p_{g}=2$ whenever $g$ is odd, while for even $g$, one has $p_{g}=3$ one-third of the time (whenever $13\mid g+$ 1). Proceeding more generally, there are $\varphi(p-1)$ primitive roots modulo the prime $p$. So by the Chinese remainder theorem, for each fixed $p$ a random $g$ satisfies $p_{g}>p$ with probability $\prod_{r\leq p}(1-\frac{\varphi(r-1)}{r})$. To make the term “probability” here rigorous, we can interpret it as limiting frequency, with $g$ sampled from integers satisfying $|g|\leq x$, where $x\to\infty$.
This probabilistic viewpoint can be used to formulate a conjecture on the upper order of $p_{g}$. While $\varphi(r-1)/r$ fluctuates as the prime $r$ varies, for the sake of estimating the above product on $r$, we can treat the terms $1-\frac{\varphi(r-1)}{r}$ as constant. More precisely, there is a certain real number $\varrho>1$ such that $\prod_{r\leq r_{k}}(1-\frac{\varphi(r-1)}{r})=\varrho^{-(1+o(1))k}$ as $k\to\infty$, where $r_{k}$ denotes the $k$th prime in the usual order. (We do not prove this here, but a related result appears as Lemma 4.1 below.) Since $12x\varrho^{-k}<$ 1 once $k>k_{0}(x):=\frac{\log{2x}}{\log{\varrho}}$, it is tempting to conjecture that $\max_{|g|\leq x}p_{g}$ is never more than about $p_{k_{0}(x)}$. (This requires “pretending” that our probabilities, which were given rigorous meaning only when fixing $k$ and sending $x$ to infinity, can be interpreted uniformly in $k$ and $x$.) This cannot be quite right, as $p_{g}=\infty$ for even square values of $g$ ! Nevertheless, it seems sensible to guess that $p_{g}\ll(\log{2|g|})(\log\log{2|g|})$ for all $g\notin\mathcal{G}$. If correct, this is sharp: In [13], Pomerance and Shparlinski report a a construction of Soundararajan yielding an infinite sequence of positive integers $g$ that (a) are all products of two distinct primes and (b) are squares modulo every odd prime $p\leq 0.7(\log{g})(\log\log{g})$. [^3] Here $0.7$ can be replaced with any constant smaller than $1/\log{4}$. These $g$ satisfy $p_{4g}\gg\log{(4g)}\log\log{(4g)}$.
This same perspective suggests that the “probability” $p_{g}>p$ is given by
$\delta_{p}:=\frac{\varphi(p-1)}{p}\prod_{r<p}\left(1-\frac{\varphi(r-1)}{r}\right).$ (6)
Taking this for granted and proceeding formally, $\mathbb{E}[p_{g}]=\sum_{p}p\delta_{p}$. Using Theorem 1.2, we give a GRH-conditional proof that this sum represents the honest average of $p_{g}$.
@@
$\lim_{x\to\infty}\frac{1}{2x}\sum_{g\in\mathcal{G},\,|g|\leq x}1=\sum_{p}p\delta_{p}.$ (7)
Here $\delta_{p}$ is as defined in (6).
(We divide by $2x$, as there are $2x+O(x^{1/2})$ integers $g\in\mathcal{G}$ with $|g|\leq x$.) There seems no hope at present of proving Corollary 1.3 unconditionally. If $p_{g}=\infty$ for even a single value of $g$, then the average becomes meaningless, and we know of no way to rule this out. Infinite values of $p_{g}$ are not the only enemy: Having $p_{g}>x\log{x}$ for some $g$, $|g|\leq x$ (along a sequence of $x$ tending to infinity) is enough to doom (7).
In an attempt to salvage the situation, one might tamp down the large values of $p_{g}$ by averaging $\min\{p_{g},\psi(x)\}$ for a threshold function $\psi$. In our final theorem, established in §5, we show that this strategy succeeds for $\psi(x)=x^{\eta}$, for any positive $\eta<\frac{1}{2}$.
@@
It would be interesting to prove Theorem 1.4 with a less stringent condition on $\eta$, such as $\eta<1$. But a substantial new idea would seem required to take $\eta$ past $1/2$. As we explain in §6, the problem becomes easier if we look instead at almost-primitive roots, meaning numbers $g$ which generate a subgroup of index at most two inside $(\mathbb{Z}/p\mathbb{Z})^{\times}$.
The problems we have taken up about $p_{g}$ are dual to those classically considered for $g_{p}$, the least primitive root modulo the prime $p$. Burgess [2] and Wang [18] have shown unconditionally that $g_{p}\ll p^{\frac{1}{4}+\varepsilon}$ for all primes $p$, while Shoup [15] (sharpening an earlier, qualitatively similar result of Wang, op. cit.) has proved under GRH that $g_{p}\ll r^{4}(1+\log{r})^{4}\log^{2}{p}$, where $r=\omega(p-1)$. Shoup’s upper bound is of size $\log^{2+o(1)}{p}$ for most primes $p$ and is always $O(\log^{6}p)$. These pointwise results are stronger than those known for $p_{g}$, but the story for average values is different. While $g_{p}$ is conjectured to have a finite, limiting mean value, this has not been established even assuming GRH (that is, the analogue of Corollary 1.3 remains open). In fact, GRH has not yielded a stronger upper bound for $\pi(x)^{-1}\sum_{p\leq x}g_{p}$ than $(\log{x})(\log\log{x})^{1+o(1)}$ (as $x\to\infty$); this last estimate is due to Elliott and Murata [3].
@@
$p\equiv 1\pmod{d}\quad\text{and}\quad g^{(p-1)/d}\equiv 1\pmod{p}$ (8)
@@
Apart from making explicit the dependence on $g$, this result is well-known and present already in [8]. Since dependence on $g$ is crucial for our purposes, we sketch a proof. We first throw out primes dividing $d|g|$; there are only $O(\log{(|g|d)})$ of these, a quantity subsumed by our error term. For the remaining primes $p$,
@@
We now turn to the proof proper. We follow Hooley’s strategy, but keep a more watchful eye on $g$-dependence in the error terms.
@@
otherwise, we say $p$ passes the $\ell$-test. Then $g$ is a primitive root modulo $p$ precisely when $p$ passes the $\ell$-test for all primes $p$. In particular, if we define
@@
For each squarefree $d\in\mathbb{N}$, let $N_{d}$ denote the count of primes $p\leq x$ which fail the $\ell$-test for each prime $\ell\mid d$. These are precisely the primes $p\leq x$ for which (8) holds, so that by Lemma 2.1 and inclusion-exclusion,
@@
$=\pi(x)\sum_{d:\,P^{+}(d)\leq\log{x}}\frac{\mu(d)}{[\mathbb{Q}(\zeta_{d},\sqrt[d]{g})]}+O\bigg{(}x^{1/2}\sum_{d:\,P^{+}(d)\leq\log{x}}\log(|g|dx)\bigg{)}.$ (9)
(Throughout, we use $P^{+}(\cdot)$ for the largest prime factor, with the convention that $P^{+}(1)=1$.) The error term is readily handled: Each squarefree $d$ with $P^{+}(d)\leq\log{x}$ satisfies $d\leq\prod_{r\leq\log{x}}r\leq x^{2}$, and there are $2^{\pi(x)}=\exp(O(\log{x}/\log\log{x}))$ such values of $d$. Hence,
$x^{1/2}\sum_{d:\,P^{+}(d)\leq\log{x}}\log(|g|dx)\ll x^{1/2}\log(|g|x)\cdot\exp(O(\log{x}/\log\log{x}))\ll x^{3/5}\log{|g|}.$ (10)
Turning to the main term, we extract from [8, pp. 213–214] that for each squarefree $d\in\mathbb{N}$,
[$\mathbb{Q}(\zeta_{d},\sqrt[d]{g}):\mathbb{Q}]=\frac{d\varphi(d)}{\varepsilon(d)\gcd(d,h)},\quad\text{where}\quad\varepsilon(d)=\begin{cases}2&\text{if $2g_{1}\mid d$ and $g_{1}\equiv 1\pmod{*}{4}$},\\ 1&\text{otherwise}.\end{cases}$
(Actually, what Hooley computes in [8] is the degree of $\mathbb{Q}(\zeta_{d},\sqrt[d_{1}]{g})$, where $d_{1}:=d/\gcd(d,h)$. But this is the same field as $\mathbb{Q}(\zeta_{d},\sqrt[d]{g})$, by Kummer theory, since $g$ and $g^{\gcd(d,h)}$ generate the same subgroup of $\mathbb{Q}(\zeta_{d})^{\times}/(\mathbb{Q}(\zeta_{d})^{\times})^{d}$.) From this, Hooley deduces in [8] that
$\sum_{d}\frac{\mu(d)}{[\mathbb{Q}(\zeta_{d}):\mathbb{Q}]}=A(g),$
where the sum is over all $d\in\mathbb{N}$. We would like to plug this result into (9), but the corresponding sum in (9) is restricted to $(\log{x})$-smooth values of $d$.
Let us examine the error incurred by replacing the sum over all $d$ by the sum over $(\log{x})$-smooth $d$. If $g_{1}\not\equiv 1\pmod{4}$, then
$\sum_{d:\,P^{+}(d)>\log{x}}\frac{\mu(d)}{[\mathbb{Q}(\zeta_{d}):\mathbb{Q}]}$ $=\sum_{d:\,P^{+}(d)>\log{x}}\mu(d)\frac{(d,h)}{d\varphi(d)}=-\sum_{\ell>\log{x}}\frac{(\ell,h)}{\ell\varphi(\ell)}\sum_{d:\,P^{+}(d)<\ell}\mu(d)\frac{(d,h)}{d\varphi(d)}$
@@
$\ll\sum_{\ell>\log{x}}\frac{(\ell,h)}{\ell(\ell-1)}\frac{\varphi(h)}{h}\prod_{\begin{subarray}{c}r\mid h\\ r\geq\ell\end{subarray}}\left(1+\frac{1}{r}\right).$
@@
$\prod_{\begin{subarray}{c}r\mid h\\ r>\log{x}\end{subarray}}\left(1+\frac{1}{r}\right)\leq\exp\bigg{(}\sum_{\begin{subarray}{c}r\mid h\\ r>\log{x}\end{subarray}}\frac{1}{r}\bigg{)}\leq\exp\bigg{(}\frac{1}{\log{x}}\sum_{\begin{subarray}{c}r\mid h\\ r>\log{x}\end{subarray}}1\bigg{)}\leq\exp\bigg{(}\frac{\log{h}}{\log x\cdot\log\log{x}}\bigg{)}\ll 1,$ (11)
@@
$\sum_{d:\,P^{+}(d)>\log{x}}\frac{\mu(d)}{[\mathbb{Q}(\zeta_{d}):\mathbb{Q}]}$ $\ll\frac{\varphi(h)}{h}\left(\sum_{\begin{subarray}{c}\ell>\log{x}\\ \ell\mid h\end{subarray}}\frac{1}{\ell}+\sum_{\begin{subarray}{c}\ell>\log{x}\\ \ell\nmid h\end{subarray}}\frac{1}{\ell^{2}}\right)$
@@
$\ll\frac{\varphi(h)}{h}\cdot\frac{\log\log{2|g|}}{\log{x}},$
where we take from the last display that $\log{h}\ll\log\log{2|g|}$.

### v2

@@ Abstract.
The celebrated Artin conjecture on primitive roots asserts that given any integer $g$ which is neither 1-1 nor a perfect square, there is an explicit constant $A(g)>0$ such that the number $\Pi(x;g)$ of primes $p\leq x$ for which $g$ is a primitive root is asymptotically $A(g)\pi(x)$ as $x\to\infty$, where $\pi(x)$ counts the number of primes not exceeding $x$. Artin’s conjecture has remained unsolved since its formulation in 1927. Nevertheless, Hooley demonstrated in 1967 that Artin’s conjecture is a consequence of the Generalized Riemann Hypothesis (GRH) for Dedekind zeta functions of certain cyclotomic-Kummer extensions over $\mathbb{Q}$. In this paper, we use GRH to establish a uniform version of the Artin–Hooley asymptotic formula. Specifically, we prove that $\Pi(x;g)\sim A(g)x/\log{x}$ whenever $\log{x}/\log\log{2|g|}\to\infty$, i.e., whenever $x$ tends to infinity faster than any power of $\log{(2|g|)}$. Under GRH, we also show that the least prime $p_{g}$ possessing $g$ as a primitive root satisfies the upper bound $p_{g}=O(\log^{19}(2|g|))$ uniformly for all non-square $g\neq-1$. We conclude with an application to the average value of $p_{g}$ and a discussion of an analogue concerning the least “almost-primitive” root.
@@ 1. Introduction
$\Pi(x;g)\sim A(g)\pi(x),\quad\text{as}\quad x\to\infty,$ (1.1)
for an explicitly given constant $A(g)>0$.
@@
$A_{0}(g)=\prod_{q\mid h}\bigg{(}1-\frac{1}{q-1}\bigg{)}\prod_{q\nmid h}\bigg{(}1-\frac{1}{q(q-1)}\bigg{)}.$ (1.2)
@@
$A_{1}(g)=1-\mu(|g_{1}|)\prod_{\begin{subarray}{c}q\mid h\\ q\mid g_{1}\end{subarray}}\frac{1}{q-2}\prod_{\begin{subarray}{c}q\nmid h\\ q\mid g_{1}\end{subarray}}\frac{1}{q^{2}-q-1};$ (1.3)
@@
It is this value of $A(g)$ for which Artin predicts the asymptotic formula (1.1). [^1] Artin’s original 1927 formulation was missing the factor of $A_{1}(g)$. Artin realized the need for $A_{1}(g)$ after learning of computations carried out by the Lehmers. See Stevenhagen’s discussion in [18].
Artin’s conjecture remains unresolved. In fact, to this day there is not a single value of $g$ for which we can show even the weaker assertion that $\Pi(x;g)\to\infty$ as $x\to\infty$. (However, work of Heath-Brown [9] implies this holds for at least one of $g=2$, 33, or $5$.) The most important progress in this direction is a 1967 theorem of Hooley [10], asserting that the full asymptotic relation (1.1) follows from the Generalized Riemann Hypothesis (GRH). [^2] Here and below, GRH means the Riemann Hypothesis for all Dedekind zeta functions of number fields.
@@
The asymptotic formula $\Pi(x;g)\sim A(g)\pi(x)$ holds whenever $\log{x}/\log\log{2|g|}\to\infty$. More precisely, there is an absolute constant $x_{0}>0$ for which the following holds: If $g\in\mathcal{G}$ and $x\geq\max\{x_{0},\log^{3}(2|g|)\}$, then
$\Pi(x;g)=A(g)\pi(x)\left(1+O\left(\frac{\log\log{x}}{\log{x}}+\frac{\log\log{2|g|}}{\log{x}}\right)\right).$ (1.4)
@@
$p_{g}\ll\log^{B}({2|g|}),$ (1.5)
for a certain absolute constant $B$. Indeed, suppose that $K$ is an admissible value of the implied constant in (1.4) and fix any constant $B>\max\{3,K\}$. If $x\geq\max\{x_{0},\log^{B}(2|g|)\}$, then
$\frac{\Pi(x;g)}{A(g)\pi(x)}$ $\geq 1-K\left(\frac{\log\log{x}}{\log{x}}+\frac{\log\log{2|g|}}{\log{x}}\right)$
$\geq 1-\frac{K}{B}-K\frac{\log\log{x}}{\log{x}}.$
The right-hand side is positive for large enough $x$, say $x\geq x_{1}=x_{1}(B,K)$, where $x_{1}$ is a constant that can be assumed to exceed $x_{0}$. It follows that $p_{g}\leq\max\{x_{1},\log^{B}(2|g|)\}$, giving (1.5).
In our next theorem, we pinpoint a numerically explicit value of $B$.
@@
The upper bound $p_{g}\ll\log^{B}(2|g|)$ holds with $B=19$.
Usually $p_{g}$ is quite small. For instance, $p_{g}=2$ whenever $g$ is odd, while for even $g$, one has $p_{g}=3$ one-third of the time (whenever $3\mid g+1$). Proceeding more generally, there are $\varphi(p-1)$ primitive roots modulo the prime $p$. So by the Chinese remainder theorem, for each fixed $p$ a random $g$ satisfies $p_{g}>p$ with probability $\prod_{r\leq p}(1-\frac{\varphi(r-1)}{r})$. To make the term “probability” here rigorous, we can interpret it as limiting frequency, with $g$ sampled from integers satisfying $|g|\leq x$, where $x\to\infty$.
This probabilistic viewpoint suggests a reasonable guess for the maximum size of $p_{g}$ when $|g|\leq x$. While $\varphi(r-1)/r$ fluctuates as the prime $r$ varies, for the sake of estimating the above product on $r$, we can treat the terms $1-\frac{\varphi(r-1)}{r}$ as constant. More precisely, there is a certain real number $\varrho>1$ such that
$\prod_{r\leq r_{k}}\left(1-\frac{\varphi(r-1)}{r}\right)=\varrho^{-(1+o(1))k}\quad\text{as}\quad k\to\infty,$
where $r_{k}$ denotes the $k$th prime in the usual order. (We prove this estimate as Lemma 5.1 below.) Hence, one might guess that for a given $k$ and $x$, the number of $g$, $|g|\leq x$, with $p_{g}>r_{k}$ is $\approx 2x\varrho^{-k}$. (Here $2x$ approximates the size of the sample space of $g$ values.) The expression $2x\varrho^{-k}$ is smaller than $1$ once $k>k_{0}(x):=\frac{\log{2x}}{\log{\varrho}}$. It is therefore tempting to conjecture that $\max_{|g|\leq x}p_{g}$ is never more than about $p_{k_{0}(x)}$. (This argument is purely heuristic; it requires “pretending” that our probabilities, which were given rigorous meaning only when fixing $k$ and sending $x$ to infinity, can be interpreted uniformly in $k$ and $x$.) This cannot be quite right, as $p_{g}=\infty$ for even square values of $g$ ! Nevertheless, it seems sensible to guess that $p_{g}\ll(\log{2|g|})(\log\log{2|g|})$ for all $g\in\mathcal{G}$. If correct, this is sharp: In [15], Pomerance and Shparlinski report a a construction of Soundararajan yielding an infinite sequence of positive integers $g$ that (a) are all products of two distinct primes and (b) are squares modulo every odd prime $p\leq 0.7(\log{g})(\log\log{g})$. [^3] Here $0.7$ can be replaced with any constant smaller than $1/\log{4}$. These $g$ satisfy $p_{4g}\gg\log{(4g)}\log\log{(4g)}$.
This same perspective suggests that the “probability” that a random integer $g$ satisfies $p_{g}=p$ is given by
$\delta_{p}:=\frac{\varphi(p-1)}{p}\prod_{r<p}\left(1-\frac{\varphi(r-1)}{r}\right).$ (1.6)
Taking this for granted and proceeding formally, $\mathbb{E}[p_{g}]=\sum_{p}p\delta_{p}$. Using Theorem 1.2, we give a GRH-conditional proof that this sum represents the honest average of $p_{g}$ over $g\in\mathcal{G}$.
@@
$\lim_{x\to\infty}\frac{1}{2x}\sum_{g\in\mathcal{G},\,|g|\leq x}p_{g}=\sum_{p}p\delta_{p}.$ (1.7)
Here $\delta_{p}$ is as defined in (1.6).
(We divide by $2x$, as there are $2x+O(x^{1/2})$ integers $g\in\mathcal{G}$ with $|g|\leq x$.) There seems no hope at present of proving Corollary 1.3 unconditionally: If $p_{g}=\infty$ for even a single value of $g\in\mathcal{G}$, then the average becomes meaningless, and we know of no way to rule this out. Infinite values of $p_{g}$ are not the only enemy: Having $p_{g}>x\log{x}$ for some $g\in\mathcal{G}$, $|g|\leq x$ (along a sequence of $x$ tending to infinity) is enough to doom (1.7).
In an attempt to salvage the situation, one might tamp down the large values of $p_{g}$ by averaging $\min\{p_{g},\psi(x)\}$ for a threshold function $\psi$. In our final theorem on $p_{g}$, established in §6, we show that this strategy succeeds for $\psi(x)=x^{\eta}$, for any positive $\eta<\frac{1}{2}$.
@@
Theorem 1.4 implies that any estimate of the shape $\max\{p_{g}:|g|\leq x,g\in\mathcal{G}\}\ll x^{\frac{1}{2}-\varepsilon}$ would suffice to establish (1.7).
It would be interesting to prove Theorem 1.4 with a less stringent condition on $\eta$, such as $\eta<1$. But a substantial new idea seems required to take $\eta$ past $1/2$. As we demonstrate in §7, the problem becomes easier if we look instead at almost-primitive roots, by which we mean the integers $g$ which generate a subgroup of index at most two inside $(\mathbb{Z}/p\mathbb{Z})^{\times}$.
The problems we have taken up about $p_{g}$ are dual to those classically considered for $g_{p}$, the least primitive root modulo the prime $p$. Burgess [2] and Wang [21] have shown unconditionally that $g_{p}\ll p^{\frac{1}{4}+\varepsilon}$ for all primes $p$, while Shoup [17] (sharpening an earlier, qualitatively similar result of Wang, op. cit.) has proved under GRH that $g_{p}\ll r^{4}(1+\log{r})^{4}\log^{2}{p}$, where $r=\omega(p-1)$. Shoup’s upper bound is of size $\log^{2+o(1)}{p}$ for most primes $p$ and is always $O(\log^{6}p)$. These pointwise results are stronger than those known for $p_{g}$, but the story for average values is different. While $g_{p}$ is conjectured to have a finite, limiting mean value as $p$ varies (among primes sampled in increasing order), this has not been established even assuming GRH. In fact, GRH has so far not yielded a stronger upper bound for $\pi(x)^{-1}\sum_{p\leq x}g_{p}$ than $(\log{x})(\log\log{x})^{1+o(1)}$ (as $x\to\infty$). This last estimate is due to Elliott and Murata [4]. In §4 of the same paper, Elliott and Murata propose a precise value for the average of $g_{p}$. Their theoretical expression is rather unwieldy and not easy to compute with. However, extensive direct computations of $g_{p}$ by Andrzej Paszkiewicz (reported on in [4]) suggest $g_{p}$ has mean value $\approx 4.924$.
2. Notation
We use standard notation for arithmetic functions throughout the paper. In particular, $\mu$ is the Möbius function, $\Lambda$ is the von Mangoldt function, $\varphi$ is the Euler totient function, and $\omega$ is the prime omega function, which, when evaluated at a nonzero integer $n$, returns the number of distinct prime factors of $n$. We write $(\cdot/\cdot)$ for the Kronecker symbol; often the “denominator” will be a prime $p$, in which case $(\cdot/p)$ may be viewed as a Legendre symbol.
Throughout, the letters $x,y,z,\delta,\varepsilon,\eta,\theta,\rho,K,Q,X,Y$ represent positive real variables, the letters $d,e,f,g,h,k,m,n,M,N$ stand for integer variables, and the letters $p,q,r,\ell$ are reserved for prime variables. The integer part of a real number $x$, which is defined as the greatest integer not exceeding $x$, will be denoted by $\lfloor x\rfloor$. We write $\gcd(m,n)$, or sometimes simply $(m,n)$, for the greatest common divisor of $m$ and $n$. For a positive integer $n$, we denote by $P^{+}(n)$ the largest prime factor of $n$, with the convention that $P^{+}(1)=1$, and by $P^{-}(n)$ the least prime factor of $n$, with the convention that $P^{-}(1)=\infty$.
We will also adopt the standard Landau–Vinogradov asymptotic notation such as $O$, $o$, $\ll$ and $\gg$, as well as the notation $\sim$ from calculus. Given real-valued functions $X,Y$ of a variable $t$ in a certain range, the relations $X=O(Y)$ and $X\ll Y$ will be used interchangeably to mean that there exists a constant $C>0$ such that $|X|\leq CY$ for all $t$ in the considered range. Next, the relation $X\gg Y$ is equivalent to $Y=O(X)$, and the relation $X=o(Y)$ is interpreted as $X/Y\to 0$ as $t\to\infty$. And as usual, we write $X\sim Y$ whenever $X/Y\to 1$ as $t\to\infty$.
[…]
@@ 2. Notation
$p\equiv 1\pmod{d}\quad\text{and}\quad g^{(p-1)/d}\equiv 1\pmod{p}$ (3.1)
@@
Apart from making explicit the dependence on $g$, this result is well-known and present already in [10]. Since dependence on $g$ is crucial for our purposes, we sketch a proof. We first throw out primes dividing $dg$; there are only $O(\log{(|g|d)})$ of these, a quantity subsumed by our error term. For the remaining primes $p$,
@@
We now turn to the proof of Theorem 1.1. We follow Hooley’s strategy, but keep a more watchful eye on $g$-dependence in the error terms.
@@
otherwise, we say $p$ passes the $\ell$-test. Then $g$ is a primitive root modulo $p$ precisely when $p$ passes the $\ell$-test for all primes $\ell$. In particular, if we define
@@
For each squarefree $d\in\mathbb{N}$, let $N_{d}$ denote the count of primes $p\leq x$ which fail the $\ell$-test for each prime $\ell\mid d$. These are precisely the primes $p\leq x$ for which (3.1) holds, so that by Lemma 3.1 and inclusion-exclusion,
@@
$=\pi(x)\sum_{d:\,P^{+}(d)\leq\log{x}}\frac{\mu(d)}{[\mathbb{Q}(\zeta_{d},\sqrt[d]{g}):\mathbb{Q}]}+O\bigg{(}x^{1/2}\sum_{d:\,P^{+}(d)\leq\log{x}}\mu(d)^{2}\log(|g|dx)\bigg{)}.$ (3.2)
The error term is readily handled: Each squarefree $d$ with $P^{+}(d)\leq\log{x}$ satisfies $d\leq\prod_{r\leq\log{x}}r\leq x^{2}$, and there are $2^{\pi(\log{x})}=\exp(O(\log{x}/\log\log{x}))$ such values of $d$. Hence,
$x^{1/2}\sum_{d:\,P^{+}(d)\leq\log{x}}\mu(d)^{2}\log(|g|dx)\ll x^{1/2}\log(|g|x)\cdot\exp(O(\log{x}/\log\log{x}))\ll x^{3/5}\log{|g|}.$ (3.3)
Turning to the main term, we extract from [10, pp. 213–214] or [20, Proposition 4.1] that for each squarefree $d\in\mathbb{N}$,
$[\mathbb{Q}(\zeta_{d},\sqrt[d]{g}):\mathbb{Q}]=\frac{d\varphi(d)}{\varepsilon(d)\gcd(d,h)},\quad\text{where}\quad\varepsilon(d)=\begin{cases}2&\text{if $2g_{1}\mid d$ and $g_{1}\equiv 1\pmod{*}{4}$},\\ 1&\text{otherwise}.\end{cases}$ (3.4)
(Actually, what Hooley computes in [10] is the degree of $\mathbb{Q}(\zeta_{d},\sqrt[d_{1}]{g})$, where $d_{1}:=d/\gcd(d,h)$. But this is the same field as $\mathbb{Q}(\zeta_{d},\sqrt[d]{g})$, by Kummer theory, since the classes of $g$ and $g^{\gcd(d,h)}$ generate the same subgroup of $\mathbb{Q}(\zeta_{d})^{\times}/(\mathbb{Q}(\zeta_{d})^{\times})^{d}$.) From this, Hooley deduces in [10] that
$\sum_{d}\frac{\mu(d)}{[\mathbb{Q}(\zeta_{d},\sqrt[d]{g}):\mathbb{Q}]}=A(g),$ (3.5)
where the sum is over all $d\in\mathbb{N}$. We would like to plug this result into (3.2), but the corresponding sum in (3.2) is restricted to $(\log{x})$-smooth values of $d$. The next lemma estimates the error incurred by replacing the sum over all $d$ by the sum over $(\log{x})$-smooth $d$.
Lemma 3.2.
We have
$\sum_{d:\,P^{+}(d)>\log{x}}\frac{\mu(d)}{[\mathbb{Q}(\zeta_{d},\sqrt[d]{g}):\mathbb{Q}]}\ll\frac{\varphi(h)}{h}\cdot\frac{\log\log{2|g|}}{\log{x}}.$ (3.6)
Proof.
If $g_{1}\not\equiv 1\pmod{4}$, then
$\sum_{d:\,P^{+}(d)>\log{x}}\frac{\mu(d)}{[\mathbb{Q}(\zeta_{d},\sqrt[d]{g}):\mathbb{Q}]}$ $=\sum_{d:\,P^{+}(d)>\log{x}}\mu(d)\frac{(d,h)}{d\varphi(d)}=-\sum_{\ell>\log{x}}\frac{(\ell,h)}{\ell\varphi(\ell)}\sum_{d:\,P^{+}(d)<\ell}\mu(d)\frac{(d,h)}{d\varphi(d)}$
@@
$\ll\sum_{\ell>\log{x}}\frac{(\ell,h)}{\ell(\ell-1)}\frac{\varphi(h)}{h}\prod_{\begin{subarray}{c}r\mid h\\ r\geq\ell\end{subarray}}\left(1+\frac{1}{r}\right).$ (3.7)
@@
$\prod_{\begin{subarray}{c}r\mid h\\ r>\log{x}\end{subarray}}\left(1+\frac{1}{r}\right)\leq\exp\bigg{(}\sum_{\begin{subarray}{c}r\mid h\\ r>\log{x}\end{subarray}}\frac{1}{r}\bigg{)}\leq\exp\bigg{(}\frac{1}{\log{x}}\sum_{\begin{subarray}{c}r\mid h\\ r>\log{x}\end{subarray}}1\bigg{)}\leq\exp\bigg{(}\frac{\log{h}}{\log x\cdot\log\log{x}}\bigg{)}\ll 1,$ (3.8)
@@
$\sum_{\ell>\log{x}}\frac{(\ell,h)}{\ell(\ell-1)}\frac{\varphi(h)}{h}\prod_{\begin{subarray}{c}r\mid h\\ r\geq\ell\end{subarray}}\left(1+\frac{1}{r}\right)$ $\ll\frac{\varphi(h)}{h}\left(\sum_{\begin{subarray}{c}\ell>\log{x}\\ \ell\mid h\end{subarray}}\frac{1}{\ell}+\sum_{\begin{subarray}{c}\ell>\log{x}\\ \ell\nmid h\end{subarray}}\frac{1}{\ell^{2}}\right)$
@@
$\ll\frac{\varphi(h)}{h}\cdot\frac{\log\log{2|g|}}{\log{x}},$ (3.10)
where we take from (3.9) that $\log{h}\ll\log\log{2|g|}$. The assertion of Lemma 3.2 now follows from (3.7) and (3.10), when $g_{1}\not\equiv 1\pmod{4}$.